OpenStudy (anonymous):

Using elimination to solve the system 2x + y = 8 and x + y = 4, which of the following could be the first step?

OpenStudy (anonymous):

A. 2x + y = 8--------> 2x + y = 8 (x + y = 4)(-1)---> -x - y = -4 B. 2x + (4 - x) = 8 C. 2x + y = x + y D. 3x + 2y = 12 E. x = 4

OpenStudy (anonymous):

so elimination... time to DRAW! :P

OpenStudy (anonymous):

What am I supposed to do?! :( Oh no... :(

OpenStudy (amistre64):

what do you do to eliminate?

OpenStudy (anonymous):